OpenStudy (anonymous):

PLEASE HELP!! A circle has its center at (2,-4) and a radius of 6 units. what is the equation of the circle?

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

We can use the standard form a a circle with center (h,k) \[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

We are given the center in form Center = (h,k) Center = (2,-4) Radius (r)= 6

OpenStudy (anonymous):

plug in the values in the equation i provided and that is the eq of the circle described

OpenStudy (anonymous):

but what would x and y be

OpenStudy (anonymous):

or would i find that out later?

OpenStudy (anonymous):

x is your input y is your output

OpenStudy (anonymous):

plug in the integers given in to you so we could see exactly what kind of circle we are talking about

OpenStudy (anonymous):

Would the answer be (X+2)^2 + (y-4)^2=6 ?

OpenStudy (anonymous):

@Deeezzzz

OpenStudy (usukidoll):

\[(x-h)^2 + (y-k)^2 = r^2 \] <- this is the equation of the circle (h,k) <- center r<- radius since we have (2,-4) with a radius 6 that means that h = 2, k = -2, and r =6 . Just plug those values into the formula

OpenStudy (anonymous):

So it would be (x-2)^2 + (y+4)^2=6^2.... Right? @Usukidoll @Deeezzzz

OpenStudy (usukidoll):

that's close. you just have to expand on the 6^2 part 6 x6 =? otherwise the equation is correct.